Divide rs 287 among 3 boys and 5 girls such that each girl gets half o what a boy gets.so each boy will get rs?
Answer by Theo(13342)

you are dividing 287 rs among 3 boys and 5 girls.
each girl gets half of what each boy gets.

let x = the amount that each girl gets.
then 2x = the amount that each boy gets.

since there are 3 boys and 5 girls, then:

3 * 2x + 5 * x = 287

simplify to get 6x + 5x = 287

combine like terms to get 11x = 287

divide both sides by 11 to get x = 287 / 11 = (26 + 1/11) rs

each girls gets (26 + 1/11) rs.

each boy gets 2 times that = (52 + 2/11) rs.

since there are 5 girls, the total amount given to the girls is 5 * (26 + 1/11) = 130 + 5/11 = (130 + 5/11) rs..

since there are 3 boys, the total amount given to the boys is 3 * (52 + 2/11) = 156 + 6/11 = (156 + 6/11) rs.

the total amount given to the boys and girls is (130 + 5/11) + (156 + 6/11) = (286 + 11/11) = 287 rs.

numbers look good.

your answer is that each girl receives (26 + 1/11) rs and each boy receives (52 + 2/11) rs.